### v1

@@ 2.2 $T\bar{T}+\Lambda_{2}$ deformation with timelike boundary
Since the dual field theory resides on the timelike boundary, the boundary Brown–York tensor is defined with Lorentzian signature [^6] Our convention is $Z_{\text{dS}}=\int\mathcal{D}\psi\,e^{-iS_{\text{dS}}}$. as
$\tilde{T}_{ab}$ $=\frac{2}{\sqrt{-h}}\frac{\delta}{\delta h^{ab}}(i\ln Z_{\text{dS}})$ (23)
@@
$\tilde{\mathcal{R}}^{(2)}-K^{2}+K_{ab}K^{ab}-2\Lambda=0.$ (24)
@@
$\tilde{T}^{a}_{a}=\frac{\ell_{\rm dS}}{16\pi G}\tilde{\mathcal{R}}^{(2)}+4\pi G\ell_{\rm dS}\left(\tilde{T}^{ab}\tilde{T}_{ab}-(\tilde{T}^{a}_{a})^{2}\right)-\frac{1}{4\pi G\ell_{\rm dS}}.$ (25)
@@
$T=\frac{\pi\lambda_{\text{dS}}}{2}\left(T^{ab}T_{ab}-(T^{a}_{a})^{2}\right)-\frac{2}{\pi\lambda_{\rm dS}}.$ (26)
@@
$\delta(i\ln Z_{\text{QFT}})=\frac{1}{2}\int d^{2}x\sqrt{-\gamma}\,T_{ab}\delta\gamma^{ab}=\frac{1}{2\lambda_{\rm dS}}\int d^{2}x\sqrt{-\gamma}\,T\,\delta\lambda_{\rm dS}.$ (28)
@@
$\partial_{\lambda_{\rm dS}}(i\ln Z_{\text{QFT}})=\frac{1}{2\lambda_{\rm dS}}\int d^{2}x\sqrt{-\gamma}\,T.$ (29)
@@
$\partial_{\lambda_{\rm dS}}S_{\text{QFT}}=\frac{\pi}{4}\int d^{2}x\,\sqrt{-\gamma}\,\left(T^{ab}T_{ab}-(T^{a}_{a})^{2}\right)-\frac{1}{\pi\lambda_{\rm dS}^{2}}\int d^{2}x\,\sqrt{-\gamma}.$ (30)
@@ 3.1 The first step: $T\bar{T}$ deformation
$\partial_{\lambda_{\rm dS}}S_{\text{QFT}}=-\frac{\pi}{2}\int dtd\phi(T_{\phi\phi}T_{tt}-T_{t\phi}^{2}).$ (33)
@@
$\partial_{\lambda_{\rm dS}}\mathcal{H}=\frac{\pi}{2}\int d\phi(T_{\phi\phi}T_{tt}-T_{t\phi}^{2}).$ (34)
@@
$=\frac{\pi}{2}\int d\phi\,\langle E,J|T_{\phi\phi}T_{tt}-T_{t\phi}^{2}|E,J\rangle$
$=\frac{\pi L}{2}\,(\langle T_{\phi\phi}\rangle\langle T_{tt}\rangle-\langle T_{t\phi}^{2}\rangle),$
where $L=\int d\phi=2\pi$ denotes the circumference of the cylinder. According to the definition of the boundary Brown–York tensor in Eq. (9), the dual stress tensor carries an additional imaginary factor $i$, and its components are therefore related to the energy, momentum, and angular momentum as AliAhmad:2025kki
$\langle T_{tt}\rangle=i\frac{E}{L},\quad\langle T_{\phi\phi}\rangle=i\frac{\partial E}{\partial L},\quad\langle T_{t\phi}\rangle=i\frac{2\pi J}{L^{2}}.$ (36)
@@
where we have taken $L=2\pi$ in the last step.
@@ 3.2 The second step: $T\bar{T}+\Lambda_{2}$ deformation
$\partial_{\lambda_{\rm dS}}S_{\text{QFT}}=\frac{\pi}{2}\int dtd\phi\left(T_{\phi\phi}T_{tt}-T_{t\phi}^{2}-\frac{2}{\pi^{2}\lambda_{{\rm dS}}^{2}}\right).$ (41)
@@
$\partial_{\lambda_{{\rm dS}}}E=\frac{\pi L}{2}\left(\langle T_{\phi\phi}\rangle\langle T_{tt}\rangle-\langle T_{t\phi}^{2}\rangle\right)-\frac{L}{\pi\lambda_{{\rm dS}}^{2}},$ (43)
@@
$\langle T_{tt}\rangle=\frac{E}{L},\quad\langle T_{\phi\phi}\rangle=-\frac{\partial E}{\partial L},\quad\langle T_{t\phi}\rangle=-\frac{2\pi J}{L^{2}}.$ (44)
@@
As shown in Fig. 3, matching the $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ flows at the critical deformation parameter $\lambda_{0}$ eliminates the complexification of the energy spectrum, allowing the deformation parameter to smoothly approach infinity. During this process, the metric of the two-dimensional manifold on which the deformed field theory resides changes from Euclidean to Lorentzian signature, reminiscent of the causal structure across the cosmological horizon in dS spacetime. In dS static coordinates, the radial coordinate $r$ is timelike outside the cosmological horizon and spacelike inside it, such that the induced metric on the $r=\text{const.}$ hypersurface possesses Euclidean signature outside the horizon and Lorentzian signature inside. This transition naturally supports the holographic interpretation of the composite deformation.
@@ 4 Bulk analysis
where $\mathcal{E}_{0}=LE_{0}=\frac{\pi\ell_{\rm dS}}{4G}$. This result precisely reproduces Eq. (38) on the field theory side with vanishing angular momentum J=0J=0101010Since the radial coordinate $r$ determines the proper circumference of the boundary, the dependence of the deformation parameter on $r$ disappears, i.e., $L=2\pi r_{c}$ and $\lambda_{{\rm dS}}=8G\ell_{\rm dS}$..

### v2

@@ 2.2 $T\bar{T}+\Lambda_{2}$ deformation with timelike boundary
Since the dual field theory resides on the timelike boundary, the Brown–York tensor is defined with Lorentzian signature [^7] Our convention is $Z_{\text{dS}}=\int\mathcal{D}\psi\,e^{-iS_{\text{dS}}}$. as
$\widetilde{T}_{(L)\,ab}$ $=\frac{2}{\sqrt{-h}}\frac{\delta}{\delta h^{ab}}(i\ln Z_{\text{dS}})$ (26)
@@
$\widetilde{\mathcal{R}}^{(2)}-K^{2}+K_{ab}K^{ab}-2\Lambda_{2}=0.$ (27)
@@
$\widetilde{T}_{(L)}=\frac{\ell_{\rm dS}}{16\pi G}\widetilde{\mathcal{R}}^{(2)}+4\pi G\ell_{\rm dS}\left(\widetilde{T}_{(L)}^{ab}\widetilde{T}_{(L)\,ab}-\widetilde{T}_{(L)}^{2}\right)-\frac{1}{4\pi G\ell_{\rm dS}}.$ (28)
@@
$T_{(L)}=\frac{\pi\lambda_{\text{dS}}}{2}\left(T_{(L)}^{ab}T_{(L)\,ab}-T_{(L)}^{2}\right)-\frac{2}{\pi\lambda_{\rm dS}}.$ (29)
@@
$\delta(i\ln Z_{\text{QFT}})=\frac{1}{2}\int d^{2}x\sqrt{-\gamma}\,T_{(L)\,ab}\delta\gamma^{ab}=\frac{1}{2\lambda_{\rm dS}}\int d^{2}x\sqrt{-\gamma}\,T_{(L)}\,\delta\lambda_{\rm dS}.$ (31)
@@
$\partial_{\lambda_{\rm dS}}(i\ln Z_{\text{QFT}})=\frac{1}{2\lambda_{\rm dS}}\int d^{2}x\sqrt{-\gamma}\,T_{(L)}.$ (32)
@@
$\partial_{\lambda_{\rm dS}}S_{\text{QFT}}=\frac{\pi}{4}\int d^{2}x\,\sqrt{-\gamma}\,\left(T_{(L)}^{ab}T_{(L)\,ab}-T_{(L)}^{2}\right)-\frac{1}{\pi\lambda_{\rm dS}^{2}}\int d^{2}x\,\sqrt{-\gamma}.$ (33)
@@ 3.1 The first step: $T\bar{T}$ deformation
$\partial_{\lambda_{\rm dS}}S_{\text{QFT}}=-\frac{\pi}{2}\int dtd\phi(T_{(E)\,\phi\phi}T_{(E)\,tt}-T_{(E)\,t\phi}^{2}).$ (36)
@@
$\partial_{\lambda_{\rm dS}}\mathcal{H}=\frac{\pi}{2}\int d\phi(T_{(E)\,\phi\phi}T_{(E)\,tt}-T_{(E)\,t\phi}^{2}).$ (37)
@@
$=\frac{\pi}{2}\int d\phi\,\langle E,J|T_{(E)\,\phi\phi}T_{(E)\,tt}-T_{(E)\,t\phi}^{2}|E,J\rangle$
$=\frac{\pi L}{2}\,(\langle T_{(E)\,\phi\phi}\rangle\langle T_{(E)\,tt}\rangle-\langle T_{(E)\,t\phi}^{2}\rangle),$
where $L=\int d\phi=2\pi$ denotes the circumference of the cylinder. According to the definition of the Brown–York tensor in Eq. (12), the dual stress tensor carries an additional imaginary factor $i$, and its components are therefore related to the energy, momentum, and angular momentum as AliAhmad:2025kki
$\langle T_{(E)\,tt}\rangle=i\frac{E}{L},\quad\langle T_{(E)\,\phi\phi}\rangle=i\frac{\partial E}{\partial L},\quad\langle T_{(E)\,t\phi}\rangle=i\frac{2\pi J}{L^{2}}.$ (39)
@@
where we have taken $L=2\pi$ in the last step. For the case of $J=0$, the critical deformation parameter is given by $\lambda_{0}=\frac{1}{E_{0}}$.
@@ 3.2 The second step: $T\bar{T}+\Lambda_{2}$ deformation
$\partial_{\lambda_{\rm dS}}S_{\text{QFT}}=\frac{\pi}{2}\int dtd\phi\left(T_{(L)\,\phi\phi}T_{(L)\,tt}-T_{(L)\,t\phi}^{2}-\frac{2}{\pi^{2}\lambda_{{\rm dS}}^{2}}\right).$ (44)
@@
$\partial_{\lambda_{{\rm dS}}}E=\frac{\pi L}{2}\left(\langle T_{(L)\,\phi\phi}\rangle\langle T_{(L)\,tt}\rangle-\langle T_{(L)\,t\phi}^{2}\rangle\right)-\frac{L}{\pi\lambda_{{\rm dS}}^{2}},$ (46)
@@
$\langle T_{(L)\,tt}\rangle=\frac{E}{L},\quad\langle T_{(L)\,\phi\phi}\rangle=-\frac{\partial E}{\partial L},\quad\langle T_{(L)\,t\phi}\rangle=-\frac{2\pi J}{L^{2}}.$ (47)
@@
As shown in Fig. 3, matching the $T\bar{T}$ and $T\bar{T}+\Lambda_{2}$ flows at the critical deformation parameter $\lambda_{0}$ avoids the complexification of the energy spectrum, allowing the deformation parameter to continuously approach infinity. We emphasize that, when the deformation parameter is increased to the critical value $\lambda_{0}$, we switch from the $T\bar{T}$ deformation to the $T\bar{T}+\Lambda_{2}$ deformation, while the signature of the manifold on which the field theory is defined changes from Euclidean to Lorentzian. This is because the $T\bar{T}$-deformed theory is defined on a spacelike surface, whereas the $T\bar{T}+\Lambda_{2}$-deformed theory is defined on a timelike surface. In the following subsection, the flows of the metric and the stress tensor will be discussed in more detail.
@@
3.3 Variational methods for solving energy spectrum
Beyond the standard approach of determining the energy spectrum via the flow equation for the action, one may also obtain the spectrum by employing a variational principle Guica:2019nzm . Furthermore, the flow of metric can be made manifest in this approach. We will take the $T\bar{T}$ deformation as an example to illustrate this method. Using the variational form (22) and the differential form (23) of the partition function, one finds
$\frac{1}{2}\partial_{\lambda}\left(\sqrt{\gamma}T_{(E)\,ab}\delta\gamma^{ab}\right)=\frac{\pi}{4}\delta\left[\sqrt{\gamma}\left(T_{(E)}^{ab}T_{(E)\,ab}-T_{(E)}^{2}\right)\right].$ (51)
Taking into account that the deformation parameters are related by $\lambda=-i\lambda_{{\rm dS}}$, and redefining the stress tensor as $T^{\prime}_{(E)\,ab}=\frac{\pi}{4}T_{(E)\,ab}$ for convenience, the above expression simplifies to
[…]
This expression can be decomposed as
$\partial_{\lambda_{{\rm dS}}}\gamma_{ab}=4i\hat{T}_{(E)\,ab},\qquad\partial_{\lambda_{{\rm dS}}}\hat{T}_{(E)\,ab}=2i\hat{T}_{(E)\,ac}\hat{T}^{\quad c}_{(E)\,b},$ (53)
where $\hat{T}_{(E)\,ab}=\left(T^{\prime}_{(E)\,ab}-\gamma_{ab}T_{(E)}^{\prime}\right)$. To avoid the computational complications arising from explicit factors of the imaginary unit $i$, we further redefine the stress tensor by setting $\ddot{T}^{\prime}_{(E)\,ab}=iT^{\prime}_{(E)\,ab}$. This leads to decoupled equations that take the same form as those obtained in Ref. AliAhmad:2025kki ,
$\partial_{\lambda_{{\rm dS}}}\gamma_{ab}=4\hat{T}^{\prime}_{(E)\,ab},\qquad\partial_{\lambda_{{\rm dS}}}\hat{T}^{\prime}_{(E)\,ab}=2\hat{T}^{\prime}_{(E)\,ac}\hat{T}^{\prime\quad c}_{(E)\,b},$ (54)
where $\hat{T}^{\prime}_{(E)\,ab}=\left(\ddot{T}^{\prime}_{(E)\,ab}-\gamma_{ab}\ddot{T}_{(E)}^{\prime}\right)$. Consequently, the metric flow is given by
$\gamma_{ab}=\gamma_{ab}^{[0]}+4\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,ab}+4\lambda_{{\rm dS}}^{2}\hat{T}^{\prime\,[0]}_{(E)\,ac}\hat{T}^{\prime\,[0]}_{(E)\,bd}\gamma^{cd\,[0]},$ (55)
while the stress tensor flow takes the form
$\hat{T}^{\prime}_{(E)\,ab}=\hat{T}^{\prime\,[0]}_{(E)\,ab}+2\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,ac}\hat{T}^{\prime\,[0]}_{(E)\,bd}\gamma^{cd\,[0]}.$ (56)
Here $\gamma_{ab}^{[0]}$ and $\hat{T}^{\prime\,[0]}_{(E)\,ab}$ denote the initial metric and the initial stress tensor at $\lambda_{{\rm dS}}=0$, respectively. In our convention, the initial metric is chosen to be the flat Euclidean metric $\gamma_{ab}^{[0]}=\delta_{ab}$. The deformed metric and the Euclidean metric are then related by a diffeomorphism,
$\gamma_{ab}$ $=\delta_{ab}+4\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,ab}+4\lambda_{{\rm dS}}^{2}\hat{T}^{\prime\,[0]}_{(E)\,ac}\hat{T}^{\prime\,[0]}_{(E)\,bd}\delta^{cd}$ (57)
$=\left(\delta_{a}^{\ c}+2\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,ae}\delta^{ec}\right)\delta_{cd}\left(\delta_{b}^{\ d}+2\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,fb}\delta^{df}\right).$
However, it should be emphasized that the $T\bar{T}$ deformation does not alter the form of the background manifold metric itself; its effect on the metric arises only through a Weyl factor that depends on the deformation parameter. In other words, the redefined stress tensor introduced here is defined solely within the coordinate system associated with the metric $\gamma_{ab}$, whereas the physical stress tensor is defined in the coordinates $y_{\lambda}$ in which the metric $\delta_{ab}$ remains Euclidean AliAhmad:2025kki . More precisely,
$\gamma_{ab}dx^{a}dx^{b}=\delta_{ab}dy^{a}_{\lambda}dy^{b}_{\lambda},\quad\frac{dy^{c}_{\lambda}}{dx^{a}}=\delta_{a}^{c}+2\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,ae}\delta^{ec}.$ (58)
The relation between the physical stress tensor $\mathcal{T}_{ab}$ and the deformed stress tensor $\ddot{T}^{\prime}_{(E)\,ab}$ is then given by
$\ddot{T}^{\prime}_{(E)\,ab}=\left(\delta_{a}^{c}+2\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,ae}\delta^{ec}\right)\mathcal{T}_{cd}\left(\delta_{b}^{d}+2\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,fb}\delta^{df}\right).$ (59)
Substituting Eqs. (56) and (57) into the above expression and simplifying, one obtains
$\ddot{T}^{\prime\,[0]}_{(E)\,ad}=\left(\delta_{a}^{c}+2\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,ae}\delta^{ec}\right)\mathcal{T}_{cd}.$ (60)
Taking into account that the undeformed physical stress tensor $\mathcal{T}_{cd}^{[0]}$ is defined in the coordinates associated with $y_{0}^{c}$, with
$\delta_{ab}dx^{a}dx^{b}=\Omega^{2}\delta_{cd}dy_{0}^{c}dy_{0}^{d},$ (61)
where $\Omega$ denotes the Weyl factor depending on the deformation parameter, one finally obtains the dependence of the physical stress tensor on the deformation parameter as
$\mathcal{T}_{cd}^{[0]}=\frac{dx^{a}}{dy_{0}^{c}}\left(\delta_{a}^{f}+2\lambda_{{\rm dS}}\hat{T}^{\prime\,[0]}_{(E)\,ae}\delta^{ef}\right)\mathcal{T}_{fb}\frac{dx^{b}}{dy_{0}^{d}}.$ (62)
In the above derivation, Eqs. (58) and (61) introduce two $\mathrm{SO}(2)$ gauge redundancies. These redundancies must be fixed by imposing the spatial periodicity of the physical coordinates $y^{a}$ together with the specified undeformed physical stress tensor. One then obtains the individual components of the physical stress tensor as
$\mathcal{T}_{tt}$ $=-\frac{1}{4\lambda_{{\rm dS}}}\left(1-\sqrt{1+8\lambda_{{\rm dS}}\mathcal{T}_{tt}^{[0]}-16\lambda_{{\rm dS}}^{2}(\mathcal{T}_{tx}^{[0]})^{2}}\right)\,$, (63)
$\mathcal{T}_{tx}$ $=\mathcal{T}_{tx}^{[0]}\,$,
$\mathcal{T}_{xx}$ $=-\frac{1-16\lambda_{{\rm dS}}^{2}(\mathcal{T}_{tx}^{[0]})^{2}-\sqrt{1+8\lambda_{{\rm dS}}\mathcal{T}_{tt}^{[0]}-16\lambda_{{\rm dS}}^{2}(\mathcal{T}_{tx}^{[0]})^{2}}}{4\lambda_{{\rm dS}}\sqrt{1+8\lambda_{{\rm dS}}\mathcal{T}_{tt}^{[0]}-16\lambda_{{\rm dS}}^{2}(\mathcal{T}_{tx}^{[0]})^{2}}}\,.$
We now return to our original convention
$\mathcal{T}_{ab}=i\frac{\pi}{4}T_{(E)\,ab}\,,\quad\mathcal{T}_{ab}^{[0]}=i\frac{\pi}{4}T_{(E)\,ab}^{[0]}\,.$ (64)
Moreover, taking into account the explicit form of the stress tensor components in the $T\bar{T}$ deformation given in Eq. (39), one finally obtains
$\small\mathcal{T}_{tt}=-\frac{\pi}{4}\frac{E}{L}=-\frac{1}{4\lambda_{{\rm dS}}}\left(1-\sqrt{1-\frac{2\pi\lambda_{{\rm dS}}E_{0}}{L}-\frac{4J^{2}\pi^{4}\lambda_{{\rm dS}}^{2}}{L^{4}}}\right).$ (65)
This result is in complete agreement with Eq. (42). In addition, by using the metric flow, one can also derive the dependence of the determinant of the metric on the deformed manifold with respect to the deformation parameter as
$\det[\gamma]$ $=\left(\det[I-2\lambda_{{\rm dS}}\delta^{-1}\hat{\mathcal{T}}_{(E)}]\right)^{-2}$ (66)
$=\left(\frac{2\sqrt{1+8\lambda_{{\rm dS}}\mathcal{T}_{tt}^{[0]}-16\lambda_{{\rm dS}}^{2}(\mathcal{T}_{tx}^{[0]})^{2}}}{1+4\lambda_{{\rm dS}}\mathcal{T}_{tx}^{[0]}+\sqrt{1+8\lambda_{{\rm dS}}\mathcal{T}_{tt}^{[0]}-16\lambda_{{\rm dS}}^{2}(\mathcal{T}_{tx}^{[0]})^{2}}}\right)^{2}$
$=\left(\frac{2\sqrt{1-\frac{2\pi\lambda_{{\rm dS}}E_{0}}{L}-\frac{4J^{2}\pi^{4}\lambda_{{\rm dS}}^{2}}{L^{4}}}}{1-\frac{\pi^{2}J\lambda_{{\rm dS}}}{2L^{2}}+\sqrt{1-\frac{2\pi\lambda_{{\rm dS}}E_{0}}{L}-\frac{4J^{2}\pi^{4}\lambda_{{\rm dS}}^{2}}{L^{4}}}}\right)^{2}\,.$
At the critical value of the deformation parameter, the determinant of the metric on the deformed manifold vanishes, indicating that the manifold becomes a null hypersurface. In Sec. 3.2, the transition of the deformation from the $T\bar{T}$ deformation to the $T\bar{T}+\Lambda_{2}$ deformation, together with the change in the signature of the manifold metric and the vanishing of the boundary metric determinant at the critical deformation parameter, is naturally reminiscent of the transition of constant-$r$ hypersurfaces in dS static coordinates from spacelike to timelike. In dS static coordinates, the radial coordinate $r$ is timelike outside the cosmological horizon and spacelike inside it. Consequently, the induced metric on the $r=\text{const.}$ hypersurface has Euclidean signature outside the horizon and Lorentzian signature inside. In particular, the cosmological horizon, being a null hypersurface, is characterized by a vanishing metric determinant. Motivated by this observation, we propose that, in dS holography, the composite flow formed by the $T\bar{T}$ flow and the $T\bar{T}+\Lambda_{2}$ flow is holographically dual to the motion of the boundary in dS static coordinates from infinity into the bulk, crossing the cosmological horizon and approaching the worldline of a static observer. This holographic proposal has the potential to unify two a priori distinct classes of dS holographic models. In the following, we test this proposal from the perspectives of quasi-local energy and holographic entanglement entropy.
@@ 4 Bulk analysis
where $\mathcal{E}_{0}=LE_{0}=\frac{\pi\ell_{\rm dS}}{4G}$. This result precisely reproduces Eq. (41) on the field theory side with vanishing angular momentum J=0J=0111111Since the radial coordinate $r$ determines the proper circumference of the boundary, the dependence of the deformation parameter on $r$ disappears, i.e., $L=2\pi r_{c}$ and $\lambda_{{\rm dS}}=8G\ell_{\rm dS}$..